Grade 12Inorganic Chemistry

Calculate crystal field stabilization energy (CFSE in terms of Dq) for the following complexes
(i) [Co(CN)6]3- (ii) [Fe(CN)6]3- (iii) [CoCl6]4-

bakhtawar
8 Years agoGrade 12

1 Answer

suvidhi mehta
8 Years ago
considering the splitting in strong field strong field andweak field ligands 
  1.   6*(-0.4)*10 Dq= -24Dq 
  2. 5*(-0.4)*10 Dq = -20Dq
  3. 5*(-0.4)*10 + 2*(0.6)*10 Dq = -8Dq